A body falls freely from the top of a tower . It covers 36% of the total height in the last second before striking the ground level. Find the height of tower?
 

Grade:11

3 Answers

Arun
6 years ago
  Let H be the height of the building and t be the time to fall through this height H. 

Then H = ½ g t^2 = 5 t^2. ----------------- (1) 

In (t-1) second the distance traveled is 1 - 0.36 H = 0.64 H. 

0.64 H = 5 (t-1)^2.---------------------------(2) 

(2) / (1) gives 0.64 = (t-1)^2/ t^2 

Or 0.8 = (t-1) / t 

Solving we get t= 5 second. 

Using (1) we get H = 125 m.
Upendra Yadav
6 years ago
 Let H be the height of the building and t be the time to fall through this height H. 

Then H = ½ g t^2 = 5 t^2. -- (1) 

In (t-1) second the distance traveled is 1 - 0.36 H = 0.64 H. 

0.64 H = 5 (t-1)^2.------(2) 

on doing  (2) / (1)we get 
    0.64 = (t-1)^2/ t^2 

Or 0.8 = (t-1) / t 

Solving we get t= 5 second. 

Using (1) we get H = 125 m
Sagar
4 years ago
Answer :- Height of the tower = 122